# Toward expert-level medical text validation with language models

### v4

@@ Abstract
With the growing use of language models (LMs) in clinical environments, there is an immediate need to evaluate the accuracy and safety of LM-generated medical text. Currently, such evaluation relies solely on manual physician review. However, detecting errors in LM-generated text is challenging because 1) manual review is costly and 2) expert-composed reference outputs are often unavailable in real-world settings. While the “LM-as-judge” paradigm (a LM evaluating another LM) offers scalable evaluation, even frontier LMs can miss subtle but clinically significant errors. To address these challenges, we propose MedVAL, a novel, self-supervised, data-efficient distillation method that leverages synthetic data to train evaluator LMs to assess whether LM-generated medical outputs are factually consistent with inputs, without requiring physician labels or reference outputs. To evaluate LM performance, we introduce MedVAL-Bench, a dataset of 840 physician-annotated outputs across 6 diverse medical tasks capturing real-world challenges, including a multilingual task reviewed by bilingual physicians. Each output is reviewed following a physician-defined taxonomy of risk levels and error categories, enabling evaluation of LMs in making deployment safety decisions. Across 10 state-of-the-art LMs spanning open-source, proprietary, and medically adapted models, MedVAL distillation significantly improves ($p<0.001$) alignment with physicians across seen and unseen tasks, increasing average F1 scores from 66% to 83%. Despite strong baseline performance, MedVAL improves the best-performing proprietary LM (GPT-4o) by 8% without training on physician-labeled data, demonstrating a performance statistically non-inferior to a single human expert ($p<0.001$). To support a scalable, risk-aware pathway towards clinical integration, we open-source: 1) Codebase, 2) MedVAL-Bench, 3) MedVAL-4B. Our benchmark provides evidence of LMs approaching expert-level ability in validating AI-generated medical text.
@@ 1 Introduction
Using a LM to evaluate another LM’s output, also termed the "LM-as-judge" paradigm, has shown promise in automating the evaluation of LM-generated text. However, most general-purpose LM-as-judge methods treat LMs as static evaluators [29], often lacking the granularity required to assess nuanced, high-stakes medical text [30, 31, 32, 33, 34]. Recent medical-specific LM-as-judge approaches suffer from certain practical limitations; these methods often assume resources like expert-labeled training data [35], availability of reference outputs [36], or retrieval-based evidence [37], none of which are readily scalable. On the other hand, many approaches focus on subdomains such as radiology (e.g., chest X-rays), limiting generalization to broader medical tasks [38, 39, 40, 41, 42, 43, 44].
@@
We evaluate MedVAL across 6 open-source and 4 proprietary LMs on MedVAL-Bench, our newly created dataset consisting of 840 physician-annotated outputs, where each output is annotated following a physician-defined taxonomy of risk levels and error categories. MedVAL-Bench includes 6 diverse tasks (focused on LM output evaluation), including medication question answering, patient query summarization, radiology report summarization, radiology impression simplification, hospital course translation, and doctor-patient dialogue summarization. To reflect the multilingual nature of U.S. clinical practice, where Spanish is the second most spoken language [46], we include an English-to-Spanish translation task annotated by bilingual physicians.
Unlike prior LM-as-judge methods that rely on zero-shot inference, MedVAL distillation improves the validation capabilities of all underlying LMs across seen and unseen tasks. MedVAL yields significant gains ($p<0.001$): average baseline F1 scores for four-class risk-grading improve from 36.7% to 51.0%, and for binary safe/unsafe classification, from 66.2% to 82.8%. Our benchmark provides evidence that LMs can achieve performance statistically non-inferior to a single human expert on risk-based validation decisions ($p<0.001$).
@@ 2.1 MedVAL Training
We propose a self-supervised distillation method for training LMs to validate LM-generated medical text, involving three stages: 1) synthetic data generation, 2) data filtering, and 3) fine-tuning. Figure 2 describes the data curation process, and Supplementary Algorithm 1 summarizes the complete training process.
@@ 2.2.1 Input Sources
The query2question task involves creating a concise query that encapsulates the essential details necessary to address the original question. We utilize the MeQSum dataset [48], which comprises: 1) health-related queries from patients, sourced from messages sent to the U.S. National Library of Medicine, and 2) corresponding abbreviated questions formulated by three medical experts to ensure that the summaries facilitate complete and accurate answer retrieval. For our study, we sample a subset of examples from both the train and test splits defined by [4], as detailed in Table 2.
@@
The report2impression task involves processing the "findings" section of a radiology exam report to create an "impression" section that succinctly highlights the critical information. We leverage the Open-i dataset [49], which consists of de-identified narrative chest x-ray reports from the Indiana Network for Patient Care database. We randomly sample train/test examples following the splits identified by [4], ensuring the number of train/test samples indicated in Table 2.
@@
The impression2simplified task involves simplifying radiology report impression sections to produce patient-friendly rewrites that preserve the clinical meaning while being understandable to a layperson. For this task, we extract the "impression" sections from the MIMIC-IV dataset [50], which contains comprehensive radiology reports sourced from patient stays in critical care units of the Beth Israel Deaconess Medical Center. Our sampling process ensured an equal distribution of imaging modalities that reflect the top imaging indications at Stanford Hospital. These include chest x-rays, CT scans of the abdomen and pelvis, CT scans of the head, MR brain studies, ultrasound exams of the pelvis, digital screening mammography, and transabdominal/transvaginal pelvic ultrasounds. Test examples were selected via stratified random sampling as indicated in Table 2. Because MIMIC-IV [50] is partially open-source, we select training examples from Open-i radiology report findings on a related but distinct report2simplified task (i.e., simplifying the findings section). For this task, we randomly sample training examples from the training split by [4], ensuring no overlap with the samples used in report2impression.
@@
The objective of dialogue2note is to summarize doctor-patient conversations into an "assessment and plan". We utilize the ACI-Bench dataset [52, 53, 54], which provides a collection of 1) 207 doctor-patient conversations and 2) patient visit notes. To ensure consistency, we randomly sample the testing examples (indicated in Table 2), from the test set defined by [4].
@@ 2.2.2 Physician Reader Study
We present the error category distribution in Supplementary Figure 9. Across 840 examples, the most common error categories are fabricated claim (45.7%), missing claim (14.0%), and incorrect recommendation (12.6%). We also observe that error frequency rises with the risk grades (for grades 1–4, average 0.14 $\to$ 3.24 errors).
@@ 2.3.1 Language Models
We evaluate a diverse collection of state-of-the-art transformer-based LMs (Supplementary Table 5). Our selection criteria included model license, context length, and size. For open-source models, we include Llama [55, 56] (Llama 3.1 8B, 3.2 3B, and 3.3 70B), Qwen3 [57] (4B dense version), and Gemma [58] (Gemma 3 27B and MedGemma 27B). For proprietary models, we include GPT-4o Mini, GPT-4o [59, 60], Claude Sonnet 4, and Gemini 2.0 Flash [61] for their enhanced performance capabilities [62].
@@ 2.3.2 Implementation Details
Unless specified, the MedVAL framework employs GPT-4o as the teacher LM, i.e., the curator of synthetic data under stages 1 and 2 from Figure 2, and filters synthetic data using $\mathcal{M}_{\text{MedVAL}}\geq\tau$, where $\tau=0.9$. This value was selected heuristically based on inspection of agreement scores to reflect a high-consistency cutoff.
@@ 3.1 Overall Performance
We show that MedVAL distillation consistently improves the performance of all LMs. The F1 score and Cohen’s $\kappa$, which reflect classification accuracy and agreement with expert-generated results, respectively, are reported in Figure 3 and Supplementary Figure 5. Across all tested LMs, MedVAL distillation leads to average baseline F1 scores improving from 36.7% to 51.0%. On average, MedVAL improved F1 scores by 113% for open-source LMs and 11% for proprietary LMs, where baseline performance was already considerably high. Despite the inherent difficulty of this four-class classification task, where even frontier LMs struggle and baseline scores never exceed 55%, MedVAL distillation improves GPT-4o’s F1 score from 54.5% to 58.7%, establishing the best performance in our benchmark.
@@ 3.3 Task-Wise Performance
The task-wise F1 scores illustrated in Table 3a (standard deviations in Supplementary Table 7) reveal substantial improvement in LM performance. On average across LMs, MedVAL distillation yielded a 65% F1 score improvement on seen tasks and an 84% improvement on unseen tasks. Notably, MedVAL displays strong improvements on dialogue2note, an unseen task with the longest input context lengths (average $\sim 1.5$k).
@@
We also conduct an in-depth analysis to understand the implications of LM predictions. For a granular understanding, we review examples where the baseline LM incorrectly evaluates the output, but corrects the evaluation after MedVAL distillation. We present such examples in Figure 4, Supplementary Figure 7, and Supplementary Figure 8 (showing "reasoning" behind LM’s risk-grade prediction). Together, these findings demonstrate that MedVAL enhances both the generalization and task-specific capabilities of diverse LMs.
@@ 3.4 Safety (Binary) Classification Performance
3.5 MedVAL Performance Ablation
Our ablation (Supplementary Figure 6) investigates the source of MedVAL’s improvements, demonstrating that gains arise from the "combined effect" of frontier LM distillation and MedVAL’s data filtering step. We evaluate two settings with no filtering ($\mathcal{M}_{\text{MedVAL}}\geq 0.0$) vs MedVAL filtering ($\mathcal{M}_{\text{MedVAL}}\geq 0.9$): 1) self-distillation (the model teaches itself), and 2) GPT-4o distillation (a frontier teacher).
@@ 3.7 Comparison with Prior Metrics
We compare MedVAL with two representative metrics from Supplementary Table 4: 1) AlignScore [31], the only metric compatible with a physician-free, reference-free, retrieval-free setup, and 2) RadGraph F1 [44], a radiology-specific metric that requires reference outputs to compute overlap. We compute Pearson correlations between each metric and physician risk grades on the report2impression subset (n=190; a radiology task containing reference outputs). As shown in Supplementary Figure 10, MedVAL correlates strongly with physicians (GPT-4o: $r=0.825$, Qwen3-4B: $r=0.833$), AlignScore is moderate ($r=0.678$), and RadGraph F1 is weak ($r=0.156$). These findings support that reference-based similarity is not a faithful proxy for reference-free, input-only clinical risk grading, whereas MedVAL aligns closely with the physician rubric.
@@ 4 Discussion
Similar to MedVAL, there exist automated methods for evaluating LM-generated medical text that have surfaced. However, these methods are either not medical domain-specific or require physicians-in-the-loop, the availability of reference outputs, or external knowledge bases for retrieval. MedVAL does not require either of these constraints. FActScore [30] uses domain-specific extractors for checking claims, though it relies on structured knowledge bases. AlignScore [31] proposes a unified alignment function for factuality assessment across tasks such as NLI, QA, and summarization. These methods, including model confidence calibration approaches [32, 33, 34], often lack the necessary nuance for clinically focused error assessment. In medical settings, there is often a continuum of right and wrong that requires a very nuanced understanding, making it uniquely challenging compared to general domains. Several studies have focused on error detection for LM-generated medical text. MedHAL [35], a benchmark for hallucination detection in medical text, relies on physician error annotations limited to specific curated medical tasks. DocLens [36] introduces multi-aspect metrics—completeness, conciseness, and attribution—tailored for medical text generation. While they report higher agreement with physician assessments than existing metrics, their evaluation relies on the availability of reference outputs. VeriFact [37] uses retrieval-based evidence to verify statements and is oriented towards multi-document summarization without the capability to train LMs. Recent studies have also focused on error assessment for text generation in radiology: ReXTrust [38] presents a fine-grained hallucination detector, ReXErr [39] injects clinical errors into radiology reports, GREEN [40] identifies significant errors in radiology reports, and FineRadScore [41] evaluates radiology report generation. While these efforts are crucial for identifying errors, their focus on radiology tasks (e.g., chest X-rays [42, 43, 44]) and reliance on reference outputs limits their generalizability. A detailed comparison of these methods is summarized in Supplementary Table 4.
@@ 5 Limitations
Despite MedVAL’s broad applicability, our study has limitations. Our prompts for output-generation and perturbation injection were manually designed, potentially yielding partly “simulated” perturbations. Our goal was to enrich the test set with error families commonly observed by physicians [76], not replace natural errors. Furthermore, about one quarter of cases are unconstrained generations (Level 1), and even in Levels 2–4 models often deviate from instructions, yielding a blend of injected and naturally occurring errors. Nevertheless, we acknowledge that broader in-the-wild evaluations remain important future work.
@@
Furthermore, a reasonable concern regarding our evaluation is that improvements could stem from simply learning the perturbation distribution. However, while supervision is synthetic (teacher signals filtered by generator-validator consistency), all evaluation uses physician labels. Moreover, gains persist on held-out tasks excluded from distillation (impression2simplified, bhc2spanish, and dialogue2note), which introduce distribution shifts. Together, these results suggest generalization beyond the perturbation style. Regardless, we leave a controlled study isolating perturbation design effects from error generalization to future works. Additionally, inter-physician agreement in our study was estimated on a stratified multi-annotated subset (15 examples $\times$ 6 tasks; 90/840) due to physician time constraints. While this design provides per-task coverage, it is a sparse sample; reliability estimates using Krippendorff’s $\alpha$ should be interpreted with this caveat. Finally, while we selected representative open-source and proprietary models for fine-tuning under practical resource settings, results may differ with larger models or other closed platforms. However, we open-source our code and trained models to allow the community to validate on broader tasks and models of interest.

### v5

@@ Abstract
With the growing use of language models (LMs) in clinical environments, there is an immediate need to evaluate the accuracy and safety of LM-generated medical text. Currently, such evaluation relies solely on manual physician review. However, detecting errors in LM-generated text is challenging because 1) manual review is costly and 2) expert-composed reference outputs are often unavailable in real-world settings. While the “LLM-as-a-judge” paradigm (a LM evaluating another LM) offers scalable evaluation, even frontier LMs can miss subtle but clinically significant errors. To address these challenges, we propose MedVAL, a novel, self-supervised, data-efficient distillation method that leverages synthetic data to train evaluator LMs to assess whether LM-generated medical outputs are factually consistent with inputs, without requiring physician labels or reference outputs. To evaluate LM performance, we introduce MedVAL-Bench, a dataset of 840 physician-annotated outputs across 6 diverse medical tasks capturing real-world challenges. Each output is reviewed following a physician-defined taxonomy of risk levels and error categories, enabling evaluation of LMs in making deployment safety decisions. Across 10 state-of-the-art LMs spanning open-source, proprietary, and medically adapted models, MedVAL distillation significantly improves ($p<0.001$) alignment with physicians across seen and unseen tasks, increasing average F1 scores from 66% to 83%. Despite strong baseline performance, MedVAL improves the best-performing proprietary LM (GPT-4o) by 8% without training on physician-labeled data, demonstrating a performance statistically non-inferior to a single human expert on a subset annotated by multiple physicians ($p<0.001$). To support a scalable, risk-aware pathway towards clinical integration, we open-source: 1) Codebase, 2) MedVAL-Bench, 3) MedVAL-4B. Our benchmark provides evidence of LMs approaching expert-level ability in validating AI-generated medical text.
@@ 1 Introduction
Using a LM to evaluate another LM’s output, also termed the "LLM-as-a-judge" paradigm, has shown promise in automating the evaluation of LM-generated text. However, most general-purpose LLM-as-a-judge methods treat LMs as static evaluators [29], often lacking the granularity required to assess nuanced, high-stakes medical text [30, 31, 32, 33, 34]. Recent medical-specific LLM-as-a-judge approaches suffer from certain practical limitations; these methods often assume resources like expert-labeled training data [35], availability of reference outputs [36], or retrieval-based evidence [37], none of which are readily scalable. On the other hand, many approaches focus on subdomains such as radiology (e.g., chest X-rays), limiting generalization to broader medical tasks [38, 39, 40, 41, 42, 43, 44].
@@
We evaluate MedVAL across 6 open-source and 4 proprietary LMs on MedVAL-Bench [46], our newly created dataset consisting of 840 physician-annotated outputs, where each output is annotated following a physician-defined taxonomy of risk levels and error categories. MedVAL-Bench includes 6 diverse tasks (focused on LM output evaluation), including medication question answering, patient query summarization, radiology report summarization, radiology impression simplification, hospital course translation, and doctor-patient dialogue summarization. To reflect the multilingual nature of U.S. clinical practice, where Spanish is the second most spoken language [47], we include an English-to-Spanish translation task annotated by bilingual physicians.
Unlike prior LLM-as-a-judge methods that rely on zero-shot inference, MedVAL improves the validation capabilities of all underlying LMs across seen and unseen tasks. MedVAL yields significant gains ($p<0.001$): average baseline F1 scores for four-class risk-grading improve from 36.7% to 51.0%, and for binary safe/unsafe classification, from 66.2% to 82.8%. Our benchmark provides evidence that LMs can achieve performance statistically non-inferior to a single human expert on a subset annotated by multiple physicians ($p<0.001$).
@@ 2.1 MedVAL Training
We propose a self-supervised distillation method for training LMs to validate LM-generated medical text, involving three stages: 1) synthetic data generation, 2) data filtering, and 3) fine-tuning. Figure 2 describes the data curation process, and Algorithm S1 summarizes the complete training process.
@@ 2.2.1 Input Sources
The query2question task involves creating a concise query that encapsulates the essential details necessary to address the original question. We utilize the MeQSum dataset [49], which comprises: 1) health-related queries from patients, sourced from messages sent to the U.S. National Library of Medicine, and 2) corresponding abbreviated questions formulated by three medical experts to ensure that the summaries facilitate complete and accurate answer retrieval. For our study, we sample a subset of examples from both the train and test splits defined by Van Veen et al. [4], as detailed in Table 2.
@@
The report2impression task involves processing the "findings" section of a radiology exam report to create an "impression" section that succinctly highlights the critical information. We leverage the Open-i dataset [50], which consists of de-identified narrative chest x-ray reports from the Indiana Network for Patient Care database. We randomly sample train/test examples following the splits identified by Van Veen et al. [4], ensuring the number of train/test samples indicated in Table 2.
@@
The impression2simplified task involves simplifying radiology report impression sections to produce patient-friendly rewrites that preserve the clinical meaning while being understandable to a layperson. For this task, we extract the "impression" sections from the MIMIC-IV dataset [51], which contains comprehensive radiology reports sourced from patient stays in critical care units of the Beth Israel Deaconess Medical Center. Our sampling process ensured an equal distribution of imaging modalities that reflect the top imaging indications at Stanford Hospital. These include chest x-rays, CT scans of the abdomen and pelvis, CT scans of the head, MR brain studies, ultrasound exams of the pelvis, digital screening mammography, and transabdominal/transvaginal pelvic ultrasounds. Test examples were selected via stratified random sampling as indicated in Table 2. Because MIMIC-IV [51] is partially open-source, we select training examples from Open-i radiology report findings on a related but distinct report2simplified task (i.e., simplifying the findings section). For this task, we randomly sample training examples from the training split by Van Veen et al. [4], ensuring no overlap with the samples used in report2impression.
@@
The objective of dialogue2note is to summarize doctor-patient conversations into an "assessment and plan". We utilize the ACI-Bench dataset [53, 54, 55], which provides a collection of 1) 207 doctor-patient conversations and 2) patient visit notes. To ensure consistency, we randomly sample the testing examples (indicated in Table 2), from the test set defined by Van Veen et al. [4].
@@ 2.2.2 Physician Reader Study
We present the risk and error category distribution in Table S3 and Figure S6. The most common error categories are fabricated claim (45.7%), missing claim (14.0%), and incorrect recommendation (12.6%). We observe that error frequency rises with the risk grades (for grades 1–4, average 0.14 $\to$ 3.24 errors).
@@ 2.3.1 Language Models
We evaluate a diverse collection of state-of-the-art transformer-based LMs (Table S2). Our selection criteria included model license, context length, and size. For open-source models, we include Llama [56, 57] (Llama 3.1 8B, 3.2 3B, and 3.3 70B), Qwen3 [58] (4B dense version), and Gemma [59] (Gemma 3 27B and MedGemma 27B). For proprietary models, we include GPT-4o Mini, GPT-4o [60, 61], Claude Sonnet 4, and Gemini 2.0 Flash [62]. The proprietary models directly serve as our state-of-the-art LLM-as-a-judge baselines [63].
@@ 2.3.2 Implementation Details
Unless specified, our framework employs GPT-4o as the curator of synthetic data under stages 1 and 2 from Figure 2, and filters synthetic data using $\mathcal{M}_{\text{MedVAL}}\geq\tau$, where $\tau=0.9$ (selected heuristically to reflect high-consistency). All results are reported using DSPy’s zero-shot chain-of-thought (CoT) prompting module.
@@ 3.1 Overall Performance
We show that MedVAL distillation consistently improves the performance of all LMs. The F1 score and Cohen’s $\kappa$, which reflect classification accuracy and agreement with expert-generated results, respectively, are reported in Figure 3 and Figure S1. Across all tested LMs, MedVAL distillation leads to average baseline F1 scores improving from 36.7% to 51.0%. On average, MedVAL improved F1 scores by 113% for open-source LMs and 11% for proprietary LMs, where baseline performance was already considerably high. Despite the inherent difficulty of this four-class classification task, where even frontier LMs struggle and baseline scores never exceed 55%, MedVAL distillation improves GPT-4o’s F1 score from 54.5% to 58.7%, establishing the best performance in our benchmark.
@@ 3.3 Task-Wise Performance
The task-wise F1 scores illustrated in Table 3a (standard deviations in Table S4) reveal substantial improvement in LM performance. On average across LMs, MedVAL distillation yielded a 65% F1 score improvement on seen tasks and an 84% improvement on unseen tasks. Notably, MedVAL displays strong improvements on dialogue2note, an unseen task with the longest input context lengths (average $\sim 1.5$k).
@@
We also conduct an in-depth analysis to understand the implications of LM predictions. For a granular understanding, we review examples where the baseline LM incorrectly evaluates the output, but corrects the evaluation after MedVAL distillation. We present such examples in Figure 4, Figure S4, and Figure S5 (showing "reasoning" behind LM’s risk-grade prediction). Together, these findings demonstrate that MedVAL enhances both the generalization and task-specific capabilities of diverse LMs.
@@ 3.4 Safety (Binary) Classification Performance
3.5 Ablations
Our ablations (Figures S2 and S3) investigate the source of MedVAL’s improvements, examining both the contribution of frontier distillation and data filtering, as well as the impact of the filtering threshold ($\tau$) on risk classification. We evaluate two settings with no filtering ($\mathcal{M}_{\text{MedVAL}}\geq 0.0$) vs MedVAL filtering ($\mathcal{M}_{\text{MedVAL}}\geq 0.9$): 1) self-distillation (the model teaches itself), and 2) GPT-4o distillation (a frontier teacher).
@@ 3.7 Comparison with Prior Metrics
We compare MedVAL with representative metrics from Table S1: 1) AlignScore [31], 2) RadGraph F1 [44], 3) BERTScore F1 [74], and 4) ROUGE-L F1 [75]. We compute Pearson correlations between each metric and physician risk grades on the report2impression subset (n=190; a radiology task containing reference outputs). As shown in Figure S7, MedVAL correlates strongly with physicians (GPT-4o: $r=0.825$, Qwen3-4B: $r=0.833$), AlignScore is moderate ($r=0.678$), while RadGraph ($r=0.156$), BERTScore ($r=0.141$), and ROUGE-L ($r=0.259$) are weak. These findings support that reference-based similarity is not a faithful proxy for reference-free, input-only clinical risk grading, whereas MedVAL aligns closely with the physician rubric.
@@ 4 Discussion
Similar to MedVAL, there exist automated methods for evaluating LM-generated medical text that have surfaced. However, these methods are either not medical domain-specific or require physicians-in-the-loop, the availability of reference outputs, or external knowledge bases for retrieval. MedVAL does not require either of these constraints. FActScore [30] uses domain-specific extractors for checking claims, though it relies on structured knowledge bases. AlignScore [31] proposes a unified alignment function for factuality assessment across tasks such as NLI, QA, and summarization. These methods, including model confidence calibration approaches [32, 33, 34], often lack the necessary nuance for clinically focused error assessment. In medical settings, there is often a continuum of right and wrong that requires a very nuanced understanding, making it uniquely challenging compared to general domains. Several studies have focused on error detection for LM-generated medical text. MedHAL [35], a benchmark for hallucination detection in medical text, relies on physician error annotations limited to specific curated medical tasks. DocLens [36] introduces multi-aspect metrics—completeness, conciseness, and attribution—tailored for medical text generation. While they report higher agreement with physician assessments than existing metrics, their evaluation relies on the availability of reference outputs. VeriFact [37] uses retrieval-based evidence to verify statements and is oriented towards multi-document summarization without the capability to train LMs. Recent studies have also focused on error assessment for text generation in radiology: ReXTrust [38] presents a fine-grained hallucination detector, ReXErr [39] injects clinical errors into radiology reports, GREEN [40] identifies significant errors in radiology reports, and FineRadScore [41] evaluates radiology report generation. While these efforts are crucial for identifying errors, their focus on radiology tasks (e.g., chest X-rays [42, 43, 44]) and reliance on reference outputs limits their generalizability. A detailed comparison of these methods is summarized in Table S1.
@@ 5 Limitations
The prompts for output-generation and perturbation were manually designed. We intended to enrich the evaluation set with error families commonly observed by physicians [79], rather than to replace naturally occurring errors. Importantly, the perturbation engine was designed with physicians-in-the-loop to reflect clinically realistic errors. Further, approximately one quarter of cases involve unconstrained generations (Level 1), and even in Levels 2-4, models frequently deviate, resulting in a mixture of injected and naturally occurring errors. Nevertheless, broader in-the-wild evaluations remain an important direction for future work.
@@
Furthermore, a reasonable concern regarding our evaluation is that improvements could stem from simply learning the perturbation distribution. However, while supervision is synthetic (teacher signals filtered by generator-validator consistency), all evaluation uses physician labels. Moreover, gains persist on held-out tasks excluded from distillation (impression2simplified, bhc2spanish, and dialogue2note), which introduce distribution shifts. Together, these results suggest generalization beyond the perturbation style. Regardless, we leave a controlled study isolating perturbation design effects from error generalization to future works. Additionally, inter-physician agreement and non-inferiority analysis in our study were estimated on a stratified multi-annotated subset (15 examples $\times$ 6 tasks; 90/840) due to physician time constraints. While this design provides per-task coverage, it is a sparse sample; reliability estimates using Krippendorff’s $\alpha$ should be interpreted with this caveat. Finally, while we selected representative open-source and proprietary models for fine-tuning under practical resource settings, results may differ with larger models or other closed platforms. However, we open-source our code and trained models to allow the community to validate on broader tasks.